OpenStudy (anonymous):

A rectangular plot, 20 ft by 18 ft, is to be used for a garden. It is decided to put a pavement inside the entire border so that 80 square feet of the plot is left for flowers. How wide should the pavement be? I'm doing the problems that I skipped when I did my online homework a while back, I tried to do this one, and I could not figure it out. One of the people in yahoo answers just gave me the wrong answers. I'm hoping you guys would be able to help me out. Thanks in advance!

OpenStudy (anonymous):

The area of the plot is \[20\times18=360ft^2\]

OpenStudy (anonymous):

Hmm it gives me [blank] feet If I just put in 360, it says its wrong and I have entered it before :(

OpenStudy (anonymous):

360-80=280 square feet will be paved..

OpenStudy (anonymous):

I have tried 280 too. It's wrong.

OpenStudy (anonymous):

wait the answer is incomplete

OpenStudy (anonymous):

I'm testing these numbers as you are sending them to me.

OpenStudy (anonymous):

you want a pavement that has the same length = 20ft solve for \[20*W = 280\]

OpenStudy (anonymous):

so W = 260?

OpenStudy (anonymous):

nope, it is \[\frac{280}{20}\]

OpenStudy (anonymous):

280/20 is equal to 14, so enter 14 ft

OpenStudy (anonymous):

nvm thats a multiplication :D not a +/-

OpenStudy (anonymous):

ugh fml...14 ft is wrong

OpenStudy (anonymous):

:-P

OpenStudy (anonymous):

Any other ideas? It could be the site that's messing up, but other students would have brought up the issue to the teacher.

OpenStudy (anonymous):

then u take the ratio of length to width which which is 18/20=.90 then lw=280 l/w=.90 where l=length and w is width

OpenStudy (anonymous):

I now know what is wrong

OpenStudy (anonymous):

solve lw=280 and l/w=.90

OpenStudy (anonymous):

try entering 4ft

OpenStudy (anonymous):

lw=(.90w)w=280, so .90w^2=280, w=17.64

OpenStudy (anonymous):

Nope 4 is wrong

OpenStudy (anonymous):

:-P

OpenStudy (anonymous):

17.64 is wrong

OpenStudy (anonymous):

wat is the answer...the problem isnt very clear

OpenStudy (anonymous):